Operations Report – January 13th

Crew 203 Operations Report 13-JAN-2019

SOL: 1

Name of person filing report: Freddy Castaneda

Non-nominal systems: Suit #9

Notes on non-nominal systems: Still pending for diagnostic.

Generator (hours run): 17hr 15min; Turned on last night (12Jan2019) at 17:29; Turned off this morning (13Jan2019) at 09:25; Turned on tonight (13Jan2019) at 17:40

Solar SOC – Turned on (12Jan2019) 83%; Turned off (13Jan2019) 100%; Turned on (13Jan2019) 85%

Diesel Reading – 55%

Propane Reading – 30%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 95%; 505 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 90%; 270 gallons

Water (loft) – At level marker 11

Static to Loft Pump used – Yes; At 16:50 to refill tank

Water Meter: 01400327

Toilet tank emptied: Yes, it was horrible 😉

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes, used for training Crew 203

Hours: 67.5

Beginning charge: 100%

Ending charge: 72%

Currently charging: Yes

Opportunity rover used: Yes, brakes are still not working correctly but it was used for training Crew 203

Hours: 45.9

Beginning charge: 100%

Ending charge: 80%

Currently charging: Yes

Curiosity rover used: Yes, used for training Crew 203

Hours: 69.3

Beginning charge: 100%

Ending charge: 88%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: Nothing to report

HabCar used and why, where? Used two times to get water in town

CrewCar used and why, where? Yes, used to get Crew 202 to town

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Suit #9 still not working; Suit #7 has not been fixed yet; I´m aware of the situation to make a diagnosis.

Summary of Hab operations: Some crew members have detected a low smell of propane when they go upstairs, we made an inspection with the propane gas detector and nothing was detected.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report